An alloy is prepared by mixing 60g copper and 70g zinc. Find the percentage by mass of copper and zinc.


1
Expert's answer
2022-06-02T17:02:03-0400

Solution:

Mass of alloy = Mass of copper + Mass of zinc = 60 g + 70 g = 130 g


Percentage by mass = (Mass of component / Mass of alloy) × 100%

Therefore,

Percentage by mass of copper = (60 g / 130 g) × 100% = 46.15%

Percentage by mass of zinc = (70 g / 130 g) × 100% = 53.85%


Answer:

The percentage by mass of copper is 46.15%

The percentage by mass of zinc is 53.85 %